OpenStudy (anonymous):

Helpppp! What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)? (x − 2)2 + (y + 3)2 = 5 (x + 2)2 + (y − 3)2 = 25 (x − 2)2 + (y + 3)2 = 25 (x − 2)2 − (y + 3)2 = 5

OpenStudy (shinalcantara):

standard equation for circle is (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center since it passes through point (-2,0) then by plotting the point you'll be able to get the radius

OpenStudy (anonymous):

(x − 2)2 + (y + 3)2 = 25

OpenStudy (shinalcantara):